OpenStudy (anonymous):

Partial Fractions(check my work): 6 ---------------------- x^3 + 4x^2 + 3x

OpenStudy (anonymous):

your work?

OpenStudy (anonymous):

\[\frac{6}{x(x^2+4x+3)}=\frac{6}{x(x+1)(x+3)}\]

OpenStudy (anonymous):

then try it